Thinking about considerations of acid/base chemistry and solubility constants. Calcium carbonate (CaCO3) is added to initially pure water that is exposed to the atmosphere with 400 ppm(v) of carbon dioxide. If solid CaCO3 remains after a long time – showing dissolution and precipitation are at equilibrium – what is the aqueous concentration of calcium ions? Use the acid constants, Henry’s Law constants and solubility products
Answer #1

The partial pressure of 400 ppm CO2 in air, considering atmospheric pressure of the mixture, is 0.3 Torr, or 3.95x10-4 Atm. Henry's constant for CO2 in water at 25°C is 29.41 amt/M, so we can calculate the concentration of CO2 in equilibrium with the air as:

3.95.c10-4atm 29.41atm/M = 1.34x10-5M IH

This dissolved CO2 is in equilibrium with the CO32- ions through two equilibriums:

CO2,(aq) + H2O1) + HCO3(aq) + Haq

and

HCO3(aq) + Co.) + H

With Ka1 and Ka2, respectively, which have values of 4.47×10−7 and 4.69×10−11. Using Ka1 we can calculate the concentration of HCO3- and H+ (which can be considered the same, since the second dissociation can be considered as negligible when compared with this one):

Ka = [H+][HC03] [CO2] CO2

These concentrations are:

[H+] = [HC0,1 = V[CO) Ka = 1.34210-5 4.47110-7 = 2.45c10-6M

And with this value we can go to the second dissociation (again, we consider the amount of H+ generated by this dissociation as negligible:

K a2= [H+][CO; [HC01

So we get:

[C0}-] = Kaz[HCO3 CO3 = -= Ka2 = 4.69c10-11 M 1+1

So this is the concentration of carbonate ion coming from the dissolution of CO2. Since we can consider air as an infinite source of CO2 at a constant pressure, this concentration is a constant.

With this, we can calculate the amount of dissolved CaCO3 using its Ksp:

Ksp = 1.4x10-8 = [Ca²+1 Coz

We already have a fixed concentration of CO32- from the dissolution of carbon dioxide, so the rest of the concentration of carbonate will come from the dissolution of CaCO3, and it will be equal to the concentration of Ca2+, since they are both coming from the same compound. If we call this concentrations x, we get:

1.4.010-8 = . (4.69.c10-11 + r) = x + 4.69210-11

This is a quadratic equation which can be solved to yield:

Ij = 1.18r10-4 M; 12 = -1.18r10-4 M

The positive answer is the only one which makes chemical sense, so the desired concentration of Ca2+ is 1.18x10-4 M.